OpenStudy (anonymous):

Find the y-intercept. y = 9x + 3 9 3

Directrix (directrix):

The y-intercept is where the graph of y = 9x + 3 crosses the y-axis. To get on the y-axis, then the x-coordinate of the point has to be 0. The y-intercept will be: y = 9*0 + 3 = 0 + 3 = 3. @Matthew1213